# Can Generated Images Serve as a Viable Modality for Text-Centric Multimodal Learning?

### v1

@@ Appendix E Appendix E: Latency and Cost Analysis
The analysis considers the three primary Text-to-Image (T2I) models used in our experiments. Costs and latencies are based on standard offerings from official APIs (OpenAI for DALL-E 3, Stability AI for Stable Diffusion models) as of late 2024. It is important to note that these values can fluctuate based on server load, API plan, and requested image quality (e.g., standard vs. HD). The following table provides a representative snapshot for generating a single image with the parameters used in our study.
@@
Model Image Size Est. Latency (sec) Est. Cost (USD)
Stable Diffusion v1.5 512x512 ~3 ~\$0.008
Stable Diffusion XL 1024x1024 ~5 ~\$0.022
DALL-E 3 (Standard) 1024x1024 ~8 ~\$0.040
The data clearly illustrates the trade-off between image quality, cost, and speed. While DALL-E 3 offers the highest semantic fidelity, it also incurs the highest cost and latency, primarily due to its integrated LLM-based prompt rewriting mechanism. Stable Diffusion XL (SDXL) presents a balanced option, while the older SD1.5 model is the fastest and cheapest, albeit at the cost of lower image quality. This quantitative analysis underscores the argument made in the main text: the selection of a T2I model for this task is a critical decision that must balance downstream performance gains against the practical constraints of a computational budget and application latency requirements.
Appendix F Appendix F: Discussion, Limitations, and Future Work

### v2

@@ Appendix E Appendix E: Latency and Cost Analysis
The analysis considers the primary Text-to-Image (T2I) models used in our experiments. Costs and latencies are based on standard offerings from official APIs (OpenAI for DALL-E 3, Stability AI for Stable Diffusion models, Black Forest Labs for Flux) as of late 2024.
@@
Model Steps Est. Latency (sec) Est. Cost (USD)
SD1.5 (Legacy) 50 ~3.0 ~\$0.008
SDXL (Standard) 50 ~5.0 ~\$0.022
SDXL-Lightning 4 ~1.2 ~\$0.006
Flux.1-schnell 4 ~0.8 ~\$0.004
DALL-E 3 N/A ~8.0 ~\$0.040
The data clearly illustrates the trade-off between image quality, cost, and speed. While DALL-E 3 offers the highest semantic fidelity, it also incurs the highest cost and latency. Flux.1-schnell represents a breakthrough, offering SDXL-level quality at a fraction of the cost and time, effectively mitigating the ”efficiency” counter-argument.
Appendix F Appendix F: Prompting Details (System Prompts)
F.1 Prompt for Image Generation (P4)
For the P4 (LLM-Elaborated Prompting) strategy, we utilized Llama-3-8B-Instruct with the following System Prompt to ensure consistent, high-quality visual descriptions:
System Prompt: ”You are an expert visual artist and photographer. Your task is to read the provided text and imagine a single, high-fidelity image that captures the core essence, entities, and atmosphere of the text. Describe this image in a detailed, comma-separated list of visual attributes, focusing on: 1. Subject (Who/What is in the center?) 2. Action/State (What is happening?) 3. Setting/Background (Where is it?) 4. Lighting/Style (e.g., ’cinematic lighting’, ’photorealistic’, ’dark and moody’). Do NOT output any conversational text. Output ONLY the visual description prompt.”
F.2 Prompt for Textual Expansion (Baseline 2)
To ensure a fair comparison, we used a similar instruction for the Textual Expansion baseline, but directed the LLM to output a natural language description to be appended to the input text, rather than a comma-separated image prompt.
System Prompt: ”You are an expert descriptive writer. Read the following text and provide a detailed, vivid visual description of the scene, objects, or atmosphere implied by the text. Your description should clarify any visual ambiguities and set the scene. Output ONLY the descriptive paragraph. Do not explain your reasoning.”
Appendix G Appendix G: Human Evaluation
To validate our automatic metrics (CLIP Score), we conducted a human evaluation on 100 randomly sampled generated images from the Amazon Reviews dataset. Three annotators rated each image on two criteria: 1. Relevance (1-5): How well does the image match the text? 2. Clarity (1-5): Is the image free of artifacts/distortions?
Table 8: Human Evaluation Results (Mean Scores).
Model Relevance Clarity
SD1.5 3.2 3.5
SDXL 4.1 4.3
Flux.1-schnell 4.0 4.5
DALL-E 3 4.6 4.4
The results align with our automatic metrics: DALL-E 3 dominates in relevance, but Flux.1-schnell is surprisingly competitive, especially in image clarity.
Appendix H Appendix H: Ablation on Fusion Architecture
We compared Concatenation (F1) vs. Cross-Attention (F2) across different learning rates to test stability.
  • Concatenation: Prone to overfitting at high learning rates (¿5e-5). Performance peaks early but degrades quickly.
  • Cross-Attention: More stable. Even at higher learning rates, the attention mechanism allows the model to ”ignore” irrelevant visual features (assigning low attention weights), acting as a soft gate. This confirms why F2 is our default choice.
Table 9: Comparison of Fusion Mechanisms across Learning Rates (Ma-F1 %).
Fusion LR=1e-5 LR=3e-5 LR=5e-5
F1 (Concat) 76.10 76.39 75.20
F2 (Cross-Attn) 76.85 77.02 76.90
Appendix I Appendix I: Visualization of Attention Maps
We visualized the cross-attention weights for the Llama-3 + Gen. Image model. Figure 5 provides a heatmap visualization of how different text tokens attend to spatial regions in the generated image.
  • Case: ”This red vacuum cleaner…”
  • Observation: The token ”red” attends strongly to the red pixels in the generated image (specifically the vacuum cleaner’s casing), while the token ”powerful” attends to the motor/body of the vacuum.
  • Conclusion: This confirms that the model is actively ”looking” at the relevant visual regions to ground the textual tokens, rather than treating the image as a generic global feature.
Figure 5: Visualization of Cross-Attention Maps. The heatmaps show the attention weight distribution of specific text tokens (e.g., ”red”, ”powerful”) over the generated image regions. Darker red/blue indicates higher attention weight.
Appendix J Appendix J: Encoder Comparison
Table 10 presents the results of using different image encoders.
Table 10: Impact of different Image Encoders on performance (Ma-F1 %).
Image Encoder Ma-F1 (%)
DINOv2 (Pure Visual) 75.80
CLIP ViT-B/32 77.02
SigLIP (SoTA Alignment) 77.45
Analysis: Although DINOv2 captures rich spatial geometric features, its lack of text alignment pre-training results in limited gains compared to SigLIP. This demonstrates that Semantic Alignment is the key factor for the success of synthetic perception.
Appendix K Appendix K: Discussion, Limitations, and Future Work